OpenStudy (anonymous):

Given: 7x + 2 = 2 + 5x – 6 Prove: x = –3 Given the equation 7x + 2 = 2 + 5x – 6, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 7x + 2 = 2 – 6 + 5x. Then, use the ____________________ to group the like terms. This gives the equation 7x + 2 = (2 – 6) + 5x. Next, combine like terms to get the equation 7x + 2 = –4 + 5x. Use the subtraction property of equality to subtract 5x from both sides of the equation. This gives the equation 2x + 2 = –4. Then use the subtraction property of equality again to subtract 2 from

OpenStudy (anonymous):

Associative property

OpenStudy (anonymous):

are you guys able to see this whole problem? i cant figure out where it looks like it cut off at in the original problem

OpenStudy (experimentx):

5x – 6 = -6+5x (comm prop*

OpenStudy (experimentx):

2-6=4, addition operation

OpenStudy (experimentx):

7x+2 = -4+5x = 5x-4 (commutative underaddition)

OpenStudy (anonymous):

ok thank-you again!

OpenStudy (experimentx):

7x+2 -5x = 5x-4 - 5x (operating both equality with same element and same operation)

OpenStudy (experimentx):

(7x+2) -5x = (5x-4) - 5x (operating both equality with same element and same operation) 7x+(2 -5x) = 5x-(4 - 5x) use commutative property, then use assocaitive property to elimniate x, repeat the same an remove 2 from lhs